# MAVIL: Design of a Multidimensional Assessment of Visual Data Literacy and its Application in a Representative Survey

### v2

@@ 5.2 Respondents’ Self-Assessed Abilities and Understanding of the Surveyed Charts
5.3 Correlative Exploration of MAVIL Variables
@@
The correlative exploration of MAVIL dimensions with scale ratings supports the creation of distinct ability dimensions: MAVIL questions within Dimension 1 through Dimension 4 show steady inner cohesion (see 4(a)) and simultaneously, these dimensions differ from one another, suggesting they query distinctive ability areas (see 4(b)). The correlative exploration highlights some variables, such as self-assessed numeracy, as significant indicators for Visual Data Literacy.
@@
Statistical analyses were conducted using Python (with statistical tests from the scipy.stats) module and R (using the MASS and psych packages). Alongside non-parametric tests, we also used parametric methods, following arguments that they are appropriate for Likert-scale data even with small sample sizes, unequal variances, or non-normal distributions [64]. The correlations (in Section 5.3.1) were calculated according to Pearson’s correlation coefficient using the median whenever results had to be aggregated over dimensions or visualizations. To cross-check and thus confirm our statistical results, we also calculated Spearman’s rank correlation coefficient for the questions within each dimension and found that the results were almost identical with deviations between the two correlation coefficients below $0.05$. For the binary results (in Section 5.3.2), a t-test for the arithmetic mean and Mood’s median test were conducted and gave similar results. Both tests gave the same result with regard to rejecting the null hypotheses. Additionally, we looked at ordinal regressions, with the outcomes supporting the hypotheses (see Figure 5).
(a) Between questions (Dim. 1–4) by visualization example.
(b) Between dimensions (Dim. 1–4) and DRS.
Figure 4: Pearson correlation tables for MAVIL. (a) shows correlations between individual questions (Dimensions 1–4), split by the visualization example shown. (b) displays correlations between MAVIL ability dimensions and DRS (based on Dimension 5), aggregated across both examples. All correlations are statistically significant ($p<0.05$). See supplementary Figures S1–S3 for detailed item-level views.
5.3.1 Correlating Assessment Variables from MAVIL
@@
This part of the correlative analysis explores relationships between MAVIL dimensions as well as between individual questions within each dimension, using statistical methods and significance testing. A consistent positive correlation of questions within Dimension 1 through Dimension 4 indicates the plausibility of constructing ability dimensions. This holds true for both Pearson and Spearman correlation coefficients. As shown in the correlation tables in 4(a) (based on Pearson), the relations between questions within dimensions for both surveyed data visualizations are fairly strong. For Dimension 1 (Comprehending) and Dimension 2 (Decoding), all coefficients fall within the moderate (0.3-0.49) to strong (0.5-0.69) range. For Dimension 3 (Aestheticizing) and Dimension 4 (Critiquing), correlations are strong to very strong (0.7-1). Note that, for Dimension 4, only the questions of Likert type (Q4a-Q4c) were considered for the correlation and that Q4d and Q4e, which were optional text fields, are evaluated separately (see Section 5.4).
@@
To evaluate internal consistency, we also calculated Cronbach’s alpha for Dimensions 1-4, which ranged from $0.75$ to $0.9$, indicating good to excellent reliability. These results further support the aggregation of items within each dimension. Given the distributional properties of the data, we opted to use the median as a robust summary measure.
@@
When dimensions were fully grouped across questions and visualizations, as done in 4(b), Dimension 1 on Comprehending strongly correlates with Dimension 2 on Decoding and Dimension 3 on Aestheticizing. Their correlation with the Data Reading Score (DRS), which corresponds to the sum of correctly solved data reading tasks per respondent, varies and remains weak (0.1-0.29) to moderate. Dimension 4 on Critiquing disrupts this dynamic, as it is the only dimension correlating negatively with Dimension 1 through Dimension 3 and the DRS. This could be due to the simplicity of the data visualizations, as they might limit the scope of finding errors (Q4a) or coming up with design suggestions (Q4c). Respondents might react differently if hidden errors or intentional design issues were to be part of the assessed visualizations.
@@ 5.3.2 Significance of Age and Numeracy
Age and numeracy correlated significantly with the number of correctly solved data reading tasks. Binary t-tests and Mood’s median test on MAVIL variables, including demographic features, were carried out in relation to the Data Reading Score (DRS). They showed that the youngest age group of the surveyed population outperformed the older respondents regarding their DRS. The age group from 18 to 24 years ($14\%$ of respondents) had a DRS median of 4 (with p=0.01). The age groups from 25 to 74 years ($86\%$ of respondents) had a DRS median equal to 3 (with p=0.01). In addition, the self-assessed numeracy level (Q6b) correlates significantly with the DRS. Those respondents who rated their numeracy level high ($37\%$ of respondents) tended to have a DRS median equal to 4 (with p=0.009). Those who rated their numeracy level neutral or low ($63\%$ of respondents) had a lower DRS median equal to 3 (with p=0.009), suggesting clear value of having respondents self-assess their ability level.
We further validated these results using ordinal regression. The outcomes showed a substantially higher probability of achieving a DRS of 4 out of 4 in the youngest age group compared to other age groups (see Figure 5(a)). Additionally, the regression confirmed the influence of self-assessed numeracy on data reading performance, as shown in the predicted probabilities in Figure 5(b).
There is no correlation between the highest level of education reported and the self-assessed level of numeracy (Pearson’s r=0.016). As expected, our analysis did not observe a meaningful difference in Visual Data Literacy dimensions based on gender.
(a) Probability for each DRS depending on the age category of the respondents. Younger respondents had a higher probability of answering all data reading tasks correctly.
(b) Probability for each DRS depending on the numeracy self-assessment of the respondents. Higher self-rated numeracy increased the likelihood of correctly answering all data reading tasks.
5.4 Report on Visualization Criticism in the Text Inputs
For each visualization, respondents had the option to provide written visualization criticism in two text fields (max. $2,000$ characters each), one for visualization strengths (Q4d) and one for visualization weaknesses (Q4e). These questions were not mandatory, yet $37\%$ submitted at least one meaningful response. All inputs for the line chart and the bar chart were categorized into i. Visualization Strengths, ii. Visualization Weaknesses and iii. Climate Commentary.
@@ 5.4.1 Categories of Text Inputs and Examples
i. Visualization Strengths were queried in Question Q4d and defined in our analysis as a category of text inputs where a clear positive point of critique related to the tested visualization was made. For the Line Chart (see 2(a)) $99$ inputs were given, such as finding the chart "inviting, easy to understand, the color blue well chosen", or that noting it was "kept quite simple not cluttered". Someone else highlighted that "the source is indicated, [which] must be highly credited". For the bar chart (see 2(b)) $113$ respective inputs were provided, describing it as "easy to understand and memorize", acknowledging that it "raises awareness", or finding "described data easily comparable".
ii. Visualization Weaknesses could be articulated for Q4e and in our analysis were defined as comments highlighting clear negative points of critique related to the tested visualization. For the line chart, $87$ inputs were provided, stating, for example, that "the graphic looks outdated, could be designed better". A few respondents were "missing axis labeling" or wondered what the uncertainty margin means. For the bar chart, there were $78$ inputs, including debate on the cloud icon, finding it "unnecessary" or "strange" (though this graph element had also been highlighted as a strength, e.g., someone said that the "cloud speaks to [them]"). Others criticized the comparison of countries, and, for instance, found the "relation to population missing" or wanted more information because "there are other countries worth mentioning that also emit high levels of CO2".
iii. Climate Commentary appeared in Q4d and Q4e and emerged in our analysis as opinions about and reactions to climate change, rather than comments on the tested visualization itself. In total, there were $35$ of such inputs. To give an impression of this category, one respondent wrote that "the so-called man-made climate change is pure fraud", others called climate change "a lie" and one respondent said that the visualization would aim towards "tricking people into fear". Others articulated feeling affected such as that the visualization "makes [them] sad" or that looking at it they felt that "the world will be destroyed".
@@ 5.4.2 Statistical Analysis of Text Inputs
A chi-squared test was employed to evaluate the relationship between the non-ordinal text inputs with demographic factors and respondents’ Data Reading Scores (DRS). For this test, i. Visualization Strengths ($212$ inputs total) and ii. Visualization Weaknesses ($165$ inputs total) were considered equally by coding them as $1$s and putting them in a 2x2 contingency table for each data visualization. Due to partially small sample sizes, significant findings from the chi-squared test were validated through an additional Monte Carlo simulation [24, 74] by randomly sampling $2,000$ replicates.
@@
For both the line and bar chart, results indicated a significant difference in text input provision across education levels (Line Chart: ${X}^{2}$=12.93, df=1, p=0.0003, Bar Chart: ${X}^{2}$=7.88, df=1, p=0.005). The simulated p-values from the Monte Carlo simulation, based on $2,000$ replicates, validate the significance of this finding (Line Chart: p=0.0005, Bar Chart: p=0.005). This suggests that respondents with secondary and tertiary educational degrees were the ones who tended to provide optional visualization criticism. As educational disparities can impact survey responses [22], possibly, higher-educated respondents may be more accustomed to survey items demanding written input and have a structural advantage in providing this type of survey answer.
@@
Moreover, a significant association was found between the provided text inputs and the DRS (Line Chart: ${X}^{2}$=10.40, df=1, p=0.001, Bar Chart: ${X}^{2}$=8.91, df=1, p=0.003). This association was validated through a Monte Carlo simulation with $2,000$ replicates (Line Chart: p=0.001, Bar Chart=0.002). These findings indicate that the respondents with a higher DRS of 3 or 4 were the ones who were more likely to articulate visualization criticism. This underscores the position that high critical ability is linked to a higher level of literacy [17]. In the chi-squared tests, no significant differences were found regarding the association of text input provision across neither age groups nor gender.
@@
For the subset of 20 respondents who provided inputs categorized as Climate Commentary (35 inputs in total), we applied a binary coding scheme (Climate Commentary: yes/ no) and conducted an independent samples t-test on the overall dataset. The results showed a mean DRS of $3.32$ for those without climate commentary and $2.9$ for those who shared an opinion, however, they were not statistically significant ($p=0.097$). While the full-sample t-test results are almost statistically significant, this appears to be driven by the imbalance in group sizes. When accounting for this imbalance through a bootstrap-like approach, the effect was not robust ($p\approx 0.23$), suggesting that the observed difference may be sensitive to sample composition rather than reflecting a consistent underlying difference.
Given the small and uneven sample sizes, respondents’ attitudes toward climate change may warrant closer examination in future research. These inputs highlight the salience of the climate discourse and that they may be valuable when analyzing respondents’ engagement with visual information, as emotional states can influence how information is extracted [46].
@@ 6 Discussion
With MAVIL, we propose a broader lens for assessing Visual Data Literacy—one that acknowledges both performance, self-perceived ability ratings, and written critique. While our goal is not to replace assessment tools like VLAT [78] or CALVI [34], MAVIL challenges established assessments by introducing additional ability dimensions and integrating underutilized measurement approaches, such as self-assessed ability ratings. This multidimensional approach may offer insights into how general audiences not only read but also perceive and reflect on their engagement and abilities related to data visualizations—insights that task-based methods alone cannot fully capture.
In this section, we reflect on the conceptual and practical limitations of our case-specific implementations of MAVIL, discuss what our survey results reveal about different dimensions of literacy, and propose directions for expanding MAVIL’s applicability. Although we focused on two simple, static climate visualizations, the results offer a compelling entry point to better understand how audiences engage with real-world data visualizations and what assumptions researchers, designers, or anyone who creates or shares visualizations might need to rethink.
6.1 Survey-Based Insights from MAVIL
@@
Traditional assessments [78, 34] excel in task-based performance, such as value retrieval or pattern identification, but overlook other dimensions of how audiences perceive and interpret visualizations. We gathered representative survey data for a German-speaking country (N=$438$) and collected insights specific to two climate data visualizations across six newly constructed Visual Data Literacy dimensions. We assessed not only performance through data reading (MAVIL Dimension 5) but also perceived comprehension (Dim1) and decoding ability (Dim2), aesthetic judgment (Dim3), critique behavior (Dim4), and context-related self-assessments like numeracy and topic interest (Dim6). While findings are specific to the visualizations used, they offer valuable implications for visualization design, information communication, and the understanding of critical data by general audiences. Even simple, static charts still pose challenges, which is exemplified by $12\%$ of respondents feeling unfamiliar with both line and bar charts (see Section 5.2).
@@
The MAVIL survey further revealed that roughly a quarter of respondents felt unable to decode data units like millimeters and millions of tons, highlighting that elements implicitly assumed to be self-explanatory by designers or curators [48] may require reconsideration. This highlights that correct interpretation by general audiences cannot be taken for granted, underscoring the value of identifying friction points in audience understanding, which can in turn inform visualization design and implementation [47]. Such findings also call into question one-size-fits-all approaches, which have been criticized for being ineffective with diverse users who have different literacy levels and types of knowledge [49, 75].
@@
Including aesthetic perception in our literacy assessment revealed that respondents who found a visualization visually appealing also tended to trust it (see Section 5.2), suggesting that, in line with existent literature [23], design and perceived credibility are closely linked. This underscores the importance of further investigating which visual elements or design choices different audience groups may perceive as appealing or trustworthy. For example, proximity techniques have been shown to increase viewer engagement [14, 46], and such insights could guide the creation of visualizations that are both inviting and credible. However, in our study, aesthetic perception had limited influence on actual interpretation, as this dimension showed only a weak correlation with data reading performance (see 4(b)).
@@
Respondents’ open-ended visualization critiques (see Section 5.4) offered valuable qualitative insight and illustrated their capacity for critical engagement. In Dimension 4, respondents provided a total of $377$ points of critique on visualizations’ strengths and weaknesses, with some comments highlighting distracting design choices or missing labels, offering grounded insight into what viewers notice and what may hinder understanding. However, in their self-ability ratings most respondents did not feel confident suggesting design improvements, indicating a gap between recognizing and articulating critique. Likely, more flawed or intentionally misleading visuals (for instance done by Camba et al. [20] or Ge et al. [34]) would elicit stronger rating responses, and future assessment implementations could explore this space.

### v3

@@ 5.2 Respondents’ Perception of the Surveyed Charts
5.3 Correlative Exploration of MdamV Variables
@@
An exploratory look at correlations between MdamV variables supports the distinction of assessment dimensions that we proposed in Sec. 3. For example, MdamV questions within Dimension 1-4 show steady inner cohesion and simultaneously, these dimensions differ from one another, suggesting they query distinctive ability areas (see Sec. 5.3.1). The correlative exploration highlights some variables, such as self-assessed numeracy, as significant indicators for visualization understanding in terms of correct data reading (see Sec. 5.3.2).
@@
Statistical analyses were conducted using Python (scipy.stats) and R (MASS, psych). Alongside non-parametric tests, we also used parametric methods, following arguments that they are appropriate for Likert-scale data even with small sample sizes, unequal variances, or non-normal distributions [70]. The correlations (in Sec. 5.3.1) were calculated according to Pearson’s correlation coefficient using the median whenever results had to be aggregated over dimensions or visualizations. To cross-check and thus confirm our statistical results, we calculated Spearman’s rank correlation coefficient for the questions within each dimension and found that the results were almost identical, with deviations between the two correlation coefficients below $0.05$. For binary results (in Sec. 5.3.2), a t-test for the arithmetic mean and Mood’s median test were conducted and gave similar results. Both tests gave the same result with regard to rejecting the null hypotheses. Additionally, we looked at ordinal regression, with the outcomes supporting the hypotheses.
Figure 4: Correlation tables for (a) MdamV Dimension 1 to 5 and (b) questions within Dimensions 1 to 4 per surveyed chart. Indicated relations are based on Pearson’s r, and all correlations are statistically significant ($p<0.05$). See supplementary Fig. S1–S3 for more detailed views.
(a) Probability for each data reading score depending on the age category of the respondents. Younger respondents had a higher probability of answering all data reading tasks correctly.
(b) Probability for each data reading score depending on the numeracy self-assessment of the respondents. Higher self-rated numeracy increased the likelihood of correctly answering all data reading tasks.
5.3.1 Correlating Assessment Variables from MdamV
@@
This part of the correlative analysis explores relationships between dimensions as well as between individual questions within dimensions, using statistical methods and significance testing. A consistent positive correlation of questions within Dimensions 1 to 4 indicates the plausibility of constructing assessment dimensions. This holds true for both Pearson and Spearman correlation coefficients. As shown in the correlation tables in Figure 4b based on Pearson’s r, the relations between questions within dimensions for both surveyed data visualizations are fairly strong. For questions on Comprehending and Decoding, all coefficients fall within the moderate (0.3-0.49) to strong (0.5-0.69) range. For questions on Aestheticizing and Critiquing, correlations are strong to very strong (0.7-1). Note that, for Critiquing, only the questions of Likert type (Q4a-c) were considered for the correlation and optional text fields are evaluated separately (see Sec. 5.4).
@@
To evaluate internal consistency, we also calculated Cronbach’s alpha for Dimensions 1 to 4, which ranged from $0.75$ to $0.9$, indicating good to excellent reliability. These results further support the aggregation of items within each dimension. Given the distributional properties of the data, we opted to use the median as a robust summary measure.
@@
When dimensions were fully grouped across questions and visualizations, as done in Figure 4a, Comprehending strongly correlates with Decoding and Aestheticizing. Their correlation with the Reading performance varies and remains weak (0.1-0.29) to moderate. Critiquing disrupts this dynamic, as it is the only dimension correlating negatively with the other ones. This could be due to the simplicity of the data visualizations, as they might limit the scope of finding errors (Q4a) or coming up with design suggestions (Q4c). Respondents might react differently if hidden errors or intentional design issues were to be part of the assessed visualizations.
@@ 5.3.2 Significance of Age and Numeracy
Age and numeracy correlated significantly with Reading performance. To test this, binary t-tests and Mood’s median test were applied to MdamV variables, including demographic features, in relation to the task-based reading questions. These responses were then aggregated into a score, summarizing the number of correct answers out of four reading questions. Testing showed that the youngest age group of the surveyed population outperformed the older respondents regarding their data reading performance. The age group from 18 to 24 years ($14\%$ of respondents) had a score median of 4 out of 4 correct answers (with p=0.01). The age groups from 25 to 74 years ($86\%$ of respondents) had a lower score median equal to 3 correct answers (with p=0.01). In addition, the self-assessed numeracy level (Q6b) correlates significantly with the data reading score. Those respondents who rated their numeracy level high ($37\%$ of respondents) tended to have a median score equal to 4 (with p=0.009). Those who rated their numeracy level neutral or low ($63\%$ of respondents) had a lower median score equal to 3 (with p=0.009). This suggests a clear value of having respondents self-assess their ability level.
We further validated these results using ordinal regression. The outcomes showed a substantially higher probability of achieving a data reading score of 4 out of 4 in the youngest age group compared to other age groups (see Fig. 5(a)). Additionally, the regression confirmed the influence of self-assessed numeracy on data reading performance, as shown in the predicted probabilities in Figure 5(b).
There is no correlation between the highest level of education reported and the self-assessed level of numeracy (Pearson’s r=0.016). Further, as expected, our analysis did not observe a meaningful difference in the provided survey answers based on gender.
5.4 Report on Visualization Critique in the Text Inputs
For each visualization, respondents had the option to provide a written visualization critique in two text fields (max. $2,000$ characters each), one for visualization strengths (Q4d) and one for visualization weaknesses (Q4e). These questions were not mandatory, yet $37\%$ submitted at least one meaningful response. All inputs for the line chart and the bar chart were categorized into i. Visualization Strengths, ii. Visualization Weaknesses and iii. Climate Commentary.
@@ 5.4.1 Categories of Text Inputs and Examples
i. Visualization Strengths were defined as a category of text inputs where a clear positive point of critique related to the tested visualization was made. For the line chart (see Fig. 2(a)) $99$ inputs were given, such as finding the chart "inviting, easy to understand, the color blue well chosen", or noting it was "kept quite simple not cluttered". Someone else highlighted that "the source is indicated, [which] must be highly credited". For the bar chart (see Fig. 2(b)) $113$ respective inputs were provided, describing it as "easy to understand and memorize", acknowledging that it "raises awareness", or finding its "data easily comparable".
ii. Visualization Weaknesses were defined as comments highlighting clear negative points of critique related to the tested visualization. For the line chart, $87$ inputs were provided, stating, for example, that "the graphic looks outdated, could be designed better". A few respondents were "missing axis labeling" or wondered what the uncertainty margin means. For the bar chart, there were $78$ inputs, including debate on the cloud icon, finding it "unnecessary" or "strange" (though this graph element had also been highlighted as a strength, e.g., someone said that the "cloud speaks to [them]"). Others criticized the comparison of countries, and, for instance, found the "relation to population missing" or wanted more information because "there are other countries worth mentioning that also emit high levels of CO2".
iii. Climate Commentary appeared in Q4d and Q4e and emerged in our analysis as opinions about and reactions to climate change, rather than comments on the tested visualization itself. In total, there were $35$ of such inputs. To give an impression of this category, one respondent wrote that "the so-called man-made climate change is pure fraud", others called climate change "a lie" and one respondent said that the visualization would aim towards "tricking people into fear". Others articulated feeling affected, such as that the visualization "makes [them] sad" or that looking at it, they felt that "the world will be destroyed".
@@ 5.4.2 Statistical Analysis of Text Inputs
A chi-squared test was employed to evaluate the relationship between the non-ordinal text inputs and demographic factors and respondents’ data reading performance, which was summarized as a score. For this test, i. Visualization Strengths ($212$ inputs total) and ii. Visualization Weaknesses ($165$ inputs total) were considered equally by coding them as $1$s and putting them in a 2x2 contingency table for each data visualization. Due to partially small sample sizes, significant findings from the chi-squared test were validated through an additional Monte Carlo simulation [25, 81] by randomly sampling $2,000$ replicates.
@@
For both the line and bar chart, results indicated a significant difference in text input provision across education levels (line chart: ${X}^{2}$=12.93, df=1, p=0.0003, bar chart: ${X}^{2}$=7.88, df=1, p=0.005). The simulated p-values, based on $2,000$ replicates, validate the significance of this finding (line chart: p=0.0005, bar chart: p=0.005). This suggests that respondents with secondary and tertiary educational degrees were the ones who tended to provide optional visualization critique. As educational disparities can impact survey responses [23], higher-educated respondents may be more accustomed to survey items demanding written input and thus have a structural propensity to provide this type of answer.
@@
Moreover, a significant association was found between the provided text inputs and the data reading performance (line chart: ${X}^{2}$=10.40, df=1, p=0.001, bar chart: ${X}^{2}$=8.91, df=1, p=0.003), and validated through a simulation with $2,000$ replicates (line chart: p=0.001, bar chart=0.002). These findings indicate that the respondents with a higher data reading score of 3 or 4 correct answers (out of 4) were the ones who were more likely to articulate a visualization critique. This underscores that critical ability can be linked to literacy levels [18]. In the chi-squared tests, no significant differences were found regarding the association of text input provision across neither age groups nor gender.
@@
For the subset of 20 respondents who provided inputs categorized as Climate Commentary (35 inputs in total), we applied a binary coding scheme (Climate Commentary: yes/ no) and conducted an independent samples t-test on the overall dataset. The results showed a mean data reading score of $3.32$ for those without climate commentary and $2.9$ for those who shared an opinion; however, they were not statistically significant ($p=0.097$). While the full-sample t-test results are almost statistically significant, this appears to be driven by the imbalance in group sizes. When accounting for this imbalance through a bootstrap-like approach, the effect was not robust ($p\approx 0.23$), suggesting that the observed difference may be sensitive to sample composition rather than reflecting a consistent underlying difference. Given the small and uneven sample sizes, respondents’ attitudes toward climate change may warrant closer examination in future research.
@@ 6 Discussion
With MdamV, we propose a broader lens for assessing visualization understanding: one that acknowledges performance, while integrating self-perceived ability ratings and voluntary written critique. Our aim is not to replace instruments like VLAT [85] or CALVI [36], but to broaden their performance-focused approach by integrating additional facets of visualization understanding into a more multidimensional method. In our survey, we applied MdamV to two climate data visualizations and gathered insights into how general audiences not only read but also perceive and reflect on their engagement with the tested data visualizations. As displayed in our results (see Sec. 5), these are insights that task-based tests alone cannot fully capture.
In this section, we reflect on the conceptual limitations of our assessment design and comment on survey-related limitations. Further, we discuss what our survey results reveal about the different assessment dimensions for visualization understanding that we proposed, and outline directions for expanding MdamV’s applicability.
6.1 Survey-Based Insights from MdamV
@@
Drawing on frameworks that deconstruct understanding into ability facets (see Sec. 2.2) and prior work related to visual data literacy (see Sec. 2.1, Sec. 3), we constructed six ability dimensions for assessing visualization understanding, bundled in a method we call MdamV. We collected insights not only on performance through data reading (MdamV Dimension 5) but also on perceived comprehension (Dim1) and decoding ability (Dim2), aesthetic judgment (Dim3), critique behavior (Dim4), and context-related self-assessments for numeracy and topic interest (Dim6). While our findings are tied to the visualizations tested, they can offer valuable insights for visualization design practices and the communication of critical data to general audiences.
@@
Our survey results show that even with some of the most common chart types [5, 67], a notable share of respondents indicated unfamiliarity with them, which is exemplified by an intersection of $12\%$ of respondents feeling unfamiliar with both line and bar charts (see Sec. 5.2). About a quarter of respondents felt unable to decode data units like millimeters and millions of tons, highlighting that elements implicitly assumed to be self-explanatory by designers or curators [52] may require reconsideration.This points out that correct interpretation by general audiences cannot be taken for granted, underscoring the value of identifying friction points in audience understanding, which can in turn inform visualization design and implementation [51]. Such findings also call into question one-size-fits-all approaches, which have been criticized for being ineffective with diverse users who have different literacy levels and types of knowledge [53, 82].
@@
Including Aestheticizing in our assessment revealed that respondents who found a visualization visually appealing also tended to trust it (see Sec. 5.2), suggesting that, in line with prior work [24], design and perceived credibility are closely linked. This underscores the value of investigating which visual elements or design choices different audience groups may perceive as appealing or trustworthy. For example, proximity techniques have been shown to increase viewer engagement [15, 50], and such insights could guide the creation of visualizations that are both inviting and credible. Our survey responses reflected this, since inputs on visualization critique pointed to reaction toward specific embellishment, such as to a $\text{CO}_{2}$ cloud icon (see Sec. 5.4). Although respondents expressed clear preferences, e.g., overall favoring the tested bar chart over the line chart (see Sec. 5.2), aesthetic perception showed only weak correlations with reading performance (see Fig. 4a), indicating that aesthetics had a limited impact on understanding.
Respondents’ open-ended critiques (see Sec. 5.4) demonstrated critical engagement and provided valuable qualitative insight into their understanding of the tested visualizations. While such inputs cannot fully capture the complexities of the reader’s critical ability, they nonetheless indicate perspective-taking. In Wiggins and McTighe’s framework [87], perspective is said to require a demonstrable shift in understanding. Although we did not test for this directly, we found that when readers step outside their role and reflect on strengths or suggest alternative design choices, they effectively adopt a designer’s viewpoint, which could represent such a shift.
@@
In Critiquing, respondents provided a total of $377$ points of critique on visualizations’ strengths and weaknesses, with some comments highlighting distracting design choices or missing labels, offering grounded insight into what viewers notice and what may hinder understanding (see Sec. 5.4). Yet, in their self-ability ratings, most respondents did not feel confident suggesting design improvements (see Sec. 5.3.1), which may indicate a gap between recognizing critique and articulating it. Possibly, more flawed or intentionally misleading visuals (for instance, done by Camba et al. [21] or Ge et al. [36]) would elicit stronger rating responses, and future applications could explore this space.